OpenStudy (anonymous):

HELP ME PLEASE DUDES AND DUDETTES .!!! The square root of negative 9 divided by (4-7i) - (6-6i)

OpenStudy (anonymous):

-6.32307692 + 6.18461538 i

OpenStudy (oaktree):

So let's tackle the numerator first. \[\sqrt{-9} = \sqrt{9} * \sqrt{-1} = 3i\]Then, the denominator is simply\[(4-7i)-(6-6i) = 4-7i-6+6i = 10-i\]So our answer is \[\frac{ 3i }{ 10-i }\]But we have to make sure that we don't have imaginary numbers in the denominator! So we get\[\frac{ 3i }{ 10-i } * \frac{ 10+i }{ 10+i } = \frac{ 30i + 3i^2 }{ 100 - 10i + 10i - i^2 } = \frac{ -3+30i }{ 101 }\] Makes sense? Let me know if you have any problems with any of the steps.

OpenStudy (oaktree):

@TakerraT, are you there?

OpenStudy (anonymous):

yes im here . thank you so much

OpenStudy (oaktree):

No problem! Is everything clear?

OpenStudy (anonymous):

i wasnt one of my answer choices... do yu wanna see my answer choices?

OpenStudy (oaktree):

Sure.

OpenStudy (oaktree):

Oh, arithmetic mistake. Second line should say \[-2-i\]So your next line would say\[\frac{ 3i }{ -2-i } * \frac{ -2+i }{ -2+i } = \frac{ -3-6i }{ 5 }\] So THAT'S your final answer. Sorry.

OpenStudy (anonymous):

a. -3-6i/5 b. 39-6i/-9 c. -39-6i/17 d. 3-6i/3

OpenStudy (anonymous):

Thank You :D

OpenStudy (oaktree):

No problem!

OpenStudy (anonymous):

(8 + 3i)(3 + 5i) :p

OpenStudy (anonymous):

wouldnt it be 24+15i^2 ?

OpenStudy (oaktree):

Yeah, but don't forget your middle terms! You have\[(8+3i)(3+5i) = 25 + 40i +9i + 15i^2 = 25 + 49i - 15 = 10+49i\]See?

OpenStudy (oaktree):

Correction: \[9+49i\]I cannot add today.

OpenStudy (anonymous):

im sooo confused lol >.<

OpenStudy (oaktree):

Are you familiar with FOIL?

OpenStudy (oaktree):

As in, do you know how to multiply out something of the form (a+b)(c+d)?

OpenStudy (anonymous):

yes i am :) im not that good in math lol

OpenStudy (oaktree):

OK. So what are you confused about?

OpenStudy (anonymous):

ill inbox yu lol >.<